Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
More Details for 2008-02-16
ISS On-Orbit Status 02/16/08

Mission 1E Flight Day 10 (FD10). Saturday - half-day off for the combined ISS and Shuttle crew except for housekeeping and voluntary work.

Wake/sleep cycle shifted back again to prepare for 2/18 undocking: 3:15am - 5:15pm EST, Shuttle crew: 3:15am - 5:45pm.

More crewtime was applied to COL (Columbus Orbital Laboratory) rack configuration, activation & operation, led by CDR Peggy Whitson and FE-2-16 Leo Eyharts. (In particular, Whitson focused on readying the EDR (European Drawer Rack), gathering equipment, outfitting the rack, installing PCDF EU (Protein Crystalization Diagnostic Facility Electronic Unit) coolant water and data connections, setting up the laptop, verifying its software load & activating it, checking out the EDR RFI (Rack Fire Indicator), and checking out the functionalities of the rack's various subsystems.)

CDR Whitson also performed Part 1 of troubleshooting & cable connectivity checkout of the PS28 (Powerstrip 28) junction box.

Afterwards Peggy relocated MSG (Microgravity Science Glovebox) payload hardware to the newly installed COL ZSR (Zero-G Stowage Rack).

MS3 Schlegel meanwhile worked on BIOLAB (BLB), installing stowage items required for BLB and BLB BioGlovebox commissioning. (BLB activation steps included installation of analysis instruments, viz., D&C (Distribution & Cleaning) cassettes for the microscope and the spectrophotometer, lamp assembly for the spectrophotometer, VTR (Video Tape Recorder) damping system, the VTR itself and its test tape, the BGB laptop (power & video cables), four BGB front filters plus removal of protective Kapton tape from BGB rear filters, EC (Experiment Container) mounting plate, and the BGB spotlight. Activities concluded with a BGB BioGlovebox containment test and sterilization functionality test.)

CDR Whitson, CDR Frick and PLT Poindexter reconfigured the O2 (oxygen) transfer system and initiated O2 pump-over from the Orbiter to the HPGTs (High-Pressure Gas Tanks) on the exterior of the U.S. Airlock (A/L). Total amount O2 available for transfer: ~90 lbs. (This transfer used the remaining operational cycles allowed under Flight Rule for the diaphragm pump of the ORCA (Oxygen Recharge Compressor Assembly) which had been modified by waiver from 280,000 to 373,000 cycles. A replacement ORCA will be delivered to KSC in December this year.)

Using the RSK1 laptop, FE-2 Malenchenko continued his test program with the new KPT-2 BAR-RM payload equipment begun on 2/13, today taking measurements with the AU-1 ultrasound analyzer at various locations in the RO (large diameter) section of the Service Module (SM) and later downlinking the data via BSR-TM channel. The data are being used for experimenting with ISS leak detection based on environmental data anomalies (temperature, humidity, and ultrasound emissions) at possible leak locations. (BAR-RM is designed to develop a procedure for detection of air leakage from ISS modules based on environmental data anomalies (temperature, humidity, ultrasound emissions). The payload uses a remote infrared thermometer (Kelvin-Video), a thermohygrometer (Iva-6A), a heat-loss anemometer/thermometer (TTM-2), an ultrasound analyzer (AU-01), and a leak detector (UT2-03) to determine physical background signs of loss of ISS pressure integrity which could be indicative of leaks in the working compartments of the station. Measurements are taken in specific zones (13 in SM PkhO and 4 in DC1), both with lights & fans turned on and off.)

FE-2 Dan Tani worked in the A/L cleaning up after yesterday's EVA-3 by Walheim and Love, deconfiguring one of the two DCS-760 cameras (without flash) and stowing it, terminating regeneration of the first set of METOX (Metal Oxide) canisters initiated by Peggy yesterday and then starting the process on the second set (#15 & #16).

CDR Whitson meanwhile conducted the weekly 10-min. CWC (Contingency Water Container) audit as part of on-going WDS (Water Delivery System) assessment of onboard water supplies. (Updated 'cue cards' based on the crew's water calldowns are sent up every other week (currently #16-0018R).)

After setting up the KUBIK-3 thermally controlled incubator in the Soyuz 15S Orbital Compartment yesterday and switching it later from +37 degC +6 degC for running overnight, Malenchenko today checked its performance at that value and later reset it to +38 degC for continued monitoring (which included removal of the resulting moisture later today).

Dan Tani removed the VOA (Volatile Organics Analyzer) from the Lab CHeCS rack and had ~2.5 hrs set aside for changing out and replacing its OMIs (On-orbit Maintenance Items) with new spares.

FE-1 Malenchenko set up the pumping equipment and initiated (later closed out) the periodic transfer of urine from four EDV-U containers in the SM to the Rodnik BV2 tank of Progress M-63/28P (emptied of potable water on 2/14).

Frick and Poindexter continued filling containers with fuel-cell generated water in the Orbiter for transfer to the ISS. After today, total water transfers amount to 11 CWCs and 8 PWRs (Payload Water Reservoirs).

The two FE-2s, Tani & Eyharts, had another hour to continue their handover activities which prepare Leo for taking over from Dan after Shuttle undock.

Servicing and maintenance tasks completed by Yuri Malenchenko today included -

The periodic (currently daily) checkout/verification of IP-1 airflow sensors in the various RS (Russian segment) hatch openings (8) in the SM, FGB and DC1 (Docking Compartment),
Completing the routine servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, plus from his discretionary 'time permitting' task list --
The daily IMS (Inventory Management System) maintenance, updating/editing its standard 'delta file' including stowage locations, for the regular weekly automated export/import to its three databases on the ground (Houston, Moscow, Baikonur),
The daily monitoring, picture-taking and downloading on the newly set up BIO-5 Rasteniya-2 ("Plants-2") experiment. (Rasteniya-2, researches growth and development of plants (peas) under spaceflight conditions in the Lada-12 greenhouse from IBMP (Institute of Bio-Medical Problems, Russian: IMBP). During its operation, the experiment requires regular daily maintenance of the experiment involving monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, and photo/video recording. During the duration of the BIO-5 experiment, students of the Moscow City Palace for Youth Creativity of the Meshchansky inter-regional center #15 in Moscow) and the Prince of Oldenburg Lyceum in St. Petersburg will be cultivating plants in parallel on the ground and conducting comparative observation of plant growth and development under gravity and zero-gravity conditions. They are receiving the photo images taken by Yuri.), and
The regular checkup on the Japanese experiment GCF-JAXA (Granada Crystallization Facility) in the Russian TBU incubator, maintained at +20 degC, including a temperature check on its ART (automatic temperature recorder). (This is a daily monitoring/temp checking, carried on the Russian voluntary task list for the duration of Expedition 16.)
A new item on the US voluntary 'job jar' task list for Peggy today was to conduct a count and audit of the contents of CD (Compact Disk) Vol. 3 container.

The crewmembers completed their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the CEVIS cycle ergometer (CDR), TVIS treadmill (FE-1, FE-2), RED resistive exercise device (CDR, FE-2) and VELO bike with bungee cord load trainer (FE-1).

Malenchenko later transferred the crew's exercise data file to the MEC (Medical Equipment Computer) for downlink, as well as the daily wristband HRM data of the workouts on RED, followed by their erasure on the HRM storage medium (done six times a week).

This morning at ~8:40am EST, the combined flight crew of 10 persons participated in a 40-min PAO TV interview/exchange from COL with a large number of media clients at NASA centers (JSC, KSC), ESA's EAC (European Astronaut Centre) in Cologne, Germany, and CNES HQ in Paris, France, for ESA. The event went exceedingly well, with every crewmember except Yuri called on to participate.

Reboost Update: This morning's reboost of the mated stack by the Orbiter was executed nominally, resulting in a burn duration 36 min 0 sec, delta-V of ~1.3 m/s and mean altitude increase ~2.3 km/1,.2 n.mi. Purpose of the reboost (the first executed by Atlantis since 2002) was to establish the correct phasing conditions for Soyuz 15S landing, Soyuz 16S launch and the next Shuttle mission, STS-123/1JA.

EVA-3 Update: Evaluation of yesterday's spacewalk showed that it ran 30 min longer than planned and achieved all objectives plus some of the get-aheads, except for 1J/A and SARJ task items.

Voluntary Science Program: For the voluntary science program on 2/19 (Tuesday), Peggy Whitson has opted for another session with the InSPACE-2 experiment.

Weekly Science Update (Expedition Sixteen -- Week 17)

ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Radiation measurements continue to be performed in the PIRS module..

ANITA: Completed.

BCAT-3 (Binary Colloidal Alloy Test 3): 'BCAT-3 has now been stowed with the intention to take it out again in about six months for the long duration sample equilibrium photos. Before BCAT-3 was stored, we were able to get some new photos of Sample 5 using the Dan Tani setup and lighting (which works really well!). With this setup, we were able to see some structure in Sample 5, which was a pleasant surprise and very exciting for the BCAT-3 Principal Investigators (PIs). Additionally, the photos that Peggy took of Dan performing the BCAT-3 experiment are 'way cool' and will be really nice to have when we are sharing the BCAT-3 results with the world.
It has been a joy to work with everyone on BCAT-3 and we're looking forward to similar opportunities with BCAT-4 (which goes up on 1JA) and with BCAT-5, which is in the works. These future experiments involve many science investigators from both NASA and CSA, which speaks volumes about how well BCAT-3 was performed by the astronauts and the ground-based support teams.'

CARDIOCOG-2: Completed.

CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): 'Dan, thanks for completing the repeat CCISS session. The session went very well, and we are very impressed with your completion of two Holter checkouts during your mission!'

CFE (Capillary Flow Experiment): Reserve.

CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.

CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.

CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): In progress.

EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.

ELITE-S2 (Elaboratore Immagini Televisive - Space 2): Planned.

EPO (Educational Payload Operations): Reserve.

ETD (Eye Tracking Device): In progress.

EuTEF (European Technology Exposure Facility): Planned. EuTEF has been successfully installed on COL External Platform during EVA-3 on 2/15 (FD9). Currently receiving telemetry on ground.

GEOFLOW: Planned.

IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Blood and urine samples of first session are stored in MELFI.

InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress..

Integrated Immune: 'Peggy, you are half way through with your mid increment Integrated Immune collections. Dan, you are half way through with your last Integrated Immune session. Leo, you are half way through with your early increment Integrated Immune session.'

KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.

LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.

MISSE (Materials ISS Experiment): Ongoing.

MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 'Pirs'.

MULTIGEN-1: MULTIGEN-1 samples will be downloaded on STS-122 (1E).

MSG-SAME (Microgravity Science Glovebox): Complete.

NOA-2 (Nitric Oxide Analyzer): Planned.


PMDIS (Perceptual Motor Deficits in Space): Complete.

SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.

SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): 'Peggy, for Sleep, your Actiwatch will stop taking data on February 20, so the team is planning to add an activity to your task list soon. As always, all your extra logging is greatly appreciated! Also, thanks for your help with CCISS this past week. Dan, we hope you enjoyed Sleeping during your mission! You will be taking off your Actiwatch on the Shuttle so we can get it on the ground and download it. The PI looks forward to all your final data!'

SOLAR (Solar Monitoring Observatory): Planned. SOLAR has been successfully installed on COL External Platform during EVA-3 on 2/15 (FD9). Currently receiving telemetry on ground.

SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.

Swab (Characterization of Microorganisms & Allergens in Spacecraft): In progress.

TRAC (Test of Reaction & Adaptation Capabilities): Planned.

ULTRASOUND: In progress.

WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): WAICO#1 hardware launched on 1E. WAICO#1 items (ATCS with chemicals, Cultivation Box Bowls with Agar) transferred to MELFI on 2/10 (FD4). Two activities for transfer of 2 x 4 WAICO#1 seed containers to MELFI were performed on 2/11to start seed vernalization.

CEO (Crew Earth Observation): 'With the extension of STS-122, and since most of us in CEO will be working late Monday night and early Tuesday morning supporting Shuttle image analysis, we will not resume CEO ops on Tuesday, 2/19 as earlier planned. We will resume CEO operations on Wednesday, Feb. 20.'

No CEO photo targets uplinked for today.

Home - Search - Browse - Alphabetic Index: 0- 1- 2- 3- 4- 5- 6- 7- 8- 9
A- B- C- D- E- F- G- H- I- J- K- L- M- N- O- P- Q- R- S- T- U- V- W- X- Y- Z
© 1997-2017 Mark Wade - Contact
© / Conditions for Use